# GaussianCity: Generative Gaussian Splatting for Unbounded 3D City Generation

### v1

@@ 1 Introduction
Extensive experiments, encompassing both qualitative and quantitative analyses on GoogleEarth [54] and KITTI-360 [28], showcase the superiority of GaussianCity over state-of-the-art methods for 3D city generation in terms of both generation quality and efficiency. The contributions are summarized as follows: 1) To the best of our knowledge, we propose the first 3D-GS generative model for unbounded 3D city generation with both high realism and efficiency. 2) We introduce BEV-Point as a highly compact representation, which ensures that the VRAM usage remains constant as the scene size scales up. 3) We present BEV-Point Decoder, which leverages Point Serializer to capture both structural and contextual characteristics of unstructured BEV points.
Figure 1: (a) Benefiting from the compact BEV-Point representation, GaussianCity can generate unbounded 3D cities using 3D Gaussian splatting (3D-GS). (b) As the number of points increases, the VRAM usage for 3D-GS grows dramatically, while the BEV-Point, serving as a compact representation, keeps the VRAM usage constant. (c) As the number of points increases, BEV-Point exhibits significantly lower growth in file storage compared to 3D-GS. (d) The proposed GaussianCity achieves not only superior generation quality but also the best efficiency in 3D city generation.
@@ 2 Related Work
Scene Generation. Unlike impressive models for generating objects and avatars, generating scenes presents greater challenges due to their extremely high diversity. Earlier approaches [26, 32] generate scenes by synthesizing videos, which lack 3D awareness and cannot ensure 3D consistency. Semantic image synthesis methods [16, 39, 48] exhibit promising results in generating scene-level content by conditioning on pixel-wise dense correspondence, such as semantic segmentation maps or depth maps. Several methods [5, 10, 59] generate natural 3D scenes by performing inpainting and outpainting on RGB images or feature maps. However, most methods can only interpolate or extrapolate a limited distance from the input views and lack generative capabilities. Recent approaches [9, 30, 54] have achieved 3D consistent scenes at an infinite scale through unbounded layout extrapolation. Another set of works [2, 12, 40] focuses on indoor scene synthesis using costly 3D datasets [11, 49] or CAD object retrieval [13, 14].
@@ 3.2 BEV-Point Initialization
In a local patch of a bird’s-eye-view (BEV) map comprising a semantic map $\mathbf{M}$ and a height field $\mathbf{H}$, a collection of BEV points within this patch can be produced by extruding the pixels in the semantic map $\mathbf{S}$ according to the corresponding values in the height field $\mathbf{H}$. We further introduce the binary density map $\mathbf{D}$ to adjust the sampling density for different semantic categories. This is driven by the observation that certain categories exhibit simpler textures (e.g., roads, water areas), allowing for reduced density to manage computational costs, while other categories (e.g., building façades) possess intricate textures, necessitating a greater number of points for representation. The coordinates of the generated BEV points, denoted as $\mathbf{C}_{F}\in\mathbb{R}^{N_{\rm pt}\times 3}$, can be generated as
@@ 3.3 BEV-Point Feature Generation
The style look-up table $\mathcal{T}$ queries the style code $\mathbf{z}_{T}^{i}$ for the instance $l$. Specifically,
@@ 4.3 Implementation Details
Hyperparameters. The grid size $g$ is set to $0.01$ and $N_{\rm PE}$ is set to $64$. The feature channels $C_{Z}$, $C_{F_{S}}$, $C_{F_{P}}$, and $C_{F_{T}}$ are set to $256$, $61$, $1280$, and $256$, respectively. The Gaussian attributes $\mathbf{A}$ consist of RGB values only (i.e., $C_{A}=3$). The loss weights $\lambda_{\rm L1}$, $\lambda_{\rm VGG}$, and $\lambda_{\rm GAN}$ are $10$, $10$, and $0.5$, respectively.
@@ 4.5 Ablation Studies
Effectiveness of BEV-Point Representation. The proposed BEV-Points, serving as a highly compact scene representation, play an important role in enabling 3D-GS to generate unbounded scenes. As shown in Figure 1(b), the VRAM usage of 3D-GS dramatically increases as the number of points increases. Typically, for a scene covering 1.5km2 with 20 million points, optimization cannot be completed on a GPU with 32GB of VRAM. In contrast, BEV-Point ensures that the number of points requiring optimization becomes a constant independent of the scene size through Ray Intersection, leading to significantly reduced VRAM usage. Besides Ray Intersection, we also evaluate the region-based and instance-based selection strategies to identify points requiring optimization. These methods can all maintain the number of points at a constant level, with “Region” indicating selection based on the camera coverage region and “Instance” indicating selection based on visible instances. After applying “Region”, “Instance”, and Ray Intersection to a scene spanning 1.5km2 with 20M points, the number of points becomes 1.4M, 680K, and 31.7K, respectively. “Region” and “Instance” are less efficient compared to the ray intersection adopted by BEV-Point, as the ray intersection maximally removes invisible points.
Effectiveness of BEV-Point Decoder. BEV-Point Decoder consists of two key components: Point Serializer and Point Transformer. As shown in Table 3, removing both components leads to a significant degeneration due to the absence of spatial correlations. Introducing only the Point Transformer partially establishes spatial correlation, but it does not achieve the same effectiveness as using both components together. Point serializer employs Equation 13 to structure the unordered BEV points. To investigate the impact of different serialization methods in point serializer, we compare the effects of using Equation 13, Hilbert curve [18] adopted in [7, 53], and both simultaneously on the generation results. As shown in Table 3, different serialization methods have minimal impact on the final generation results. For efficiency reasons, we use Equation 13 in Point Serializer due to its lower computational complexity. Refer to Figures 6 and 7 in the appendix for more qualitative comparisons.
4.6 Limitations
While our method demonstrates promising results for unconditional 3D city generation, it still has several limitations. Firstly, the BEV-Point Initialization elevates points to a height based on the corresponding value in the height field, which assumes buildings adhere to the Manhattan assumption and cannot represent hollow structures. Secondly, the BEV-Point decoder does not fully exploit the expressive capacity of 3D-GS, as it only predicts RGB in the Gaussian attributes. However, predicting an excessive number of attributes simultaneously may destabilize the training process.
Table 2: Effectiveness of BEV-Point Decoder. Note that “Ser.” and “Trans” denote “Point Serializer” and “Point Transformer”, respectively.
Ser. Trans. FID $\downarrow$ KID $\downarrow$ DE $\downarrow$ CE $\downarrow$
@@ 4.6 Limitations
Table 3: Effectiveness of different serialization methods. Note that “Eq. 13” denote the serialization used in Point Serializer.
@@ 5 Conclusion
In this paper, we introduce GaussianCity for generating unbounded 3D cities. It is the first method to employ 3D Gaussian splatting (3D-GS) for unbounded scene generation. We introduce BEV-Point, a highly compact scene representation to mitigate the substantial increase in VRAM usage associated with representing large-scale scenes using 3D-GS. Additionally, we present BEV-Point Decoder, which leverages Point Serializer to capture the structural and contextual characteristics of unordered BEV points. Compared to previous NeRF-based city generators, our approach achieves better visual results and a 60-fold increase in runtime speed compared to CityDreamer. Extensive experimental results on GoogleEarth and KITTI-360 demonstrate the effectiveness of the proposed GaussianCity in generating both drone-view and street-view in city scenes.
@@ A.2 Discussion on Generated Gaussian Attributes
Due to the carefully designed BEV-Point initialization, all points are evenly distributed on the surface of objects. Therefore, Gaussian attributes other than RGB can be left unestimated and defaulted. Table 4 and Figure 8 provide a quantitative and qualitative comparison of various attributes generated for 3D Gaussians, demonstrating that optimizing 3D Gaussian attributes beyond RGB not only complicates network convergence but also significantly impacts the quality of the generated results.
@@ B.1 Building Interpolation
As shown in Figure 9, GaussianCity showcases the capability to interpolate building styles controlled by the variable $z$.
Figure 9: Linear interpolation along the building style. The style of each building gradually changes from left to right. In the first row, only the styles of the highlighted buildings are altered, while in the second row, the styles of all buildings are changed.
@@ B.3 Additional Qualitative Comparison
In Figures 11 and 12, we provide more visual comparisons with state-of-the-art methods on GoogleEarth and KITTI-360, respectively.

### v2

# Generative Gaussian Splatting for Unbounded 3D City Generation

@@ Abstract
Figure 1: (a) Benefiting from the compact BEV-Point representation, GaussianCity can generate unbounded 3D cities using 3D Gaussian splatting (3D-GS). (b) As the number of points increases, VRAM usage during 3D-GS training rises significantly, whereas BEV-Point, acting as a compact representation, maintains a constant VRAM usage. (c) As the number of points increases, BEV-Point exhibits significantly lower growth in file storage compared to 3D-GS. (d) The proposed GaussianCity achieves not only superior generation quality but also the best efficiency in 3D city generation.
@@ 1 Introduction
The contributions are summarized as follows:
  • To the best of our knowledge, we propose the first 3D-GS generative model for unbounded 3D city generation with both high realism and efficiency.
  • We introduce BEV-Point as a highly compact representation, which ensures that the VRAM usage remains constant as the scene size scales up.
  • We present BEV-Point Decoder, which leverages Point Serializer to capture both structural and contextual characteristics of unstructured BEV points.
  • Extensive experiments showcase the superiority of GaussianCity over state-of-the-art methods for 3D city generation in terms of both generation quality and efficiency.
@@ 2 Related Work
Scene Generation. Unlike impressive models for generating objects and avatars, generating scenes presents greater challenges due to their extremely high diversity. Earlier approaches [27, 33] generate scenes by synthesizing videos, which lack 3D awareness and cannot ensure 3D consistency. Semantic image synthesis methods [17, 40, 49] exhibit promising results in generating scene-level content by conditioning on pixel-wise dense correspondence, such as semantic segmentation maps or depth maps. Several methods [6, 11, 59] generate natural 3D scenes by performing inpainting and outpainting on RGB images or feature maps. However, most methods can only interpolate or extrapolate a limited distance from the input views and lack generative capabilities. Recent approaches [10, 31, 55] have achieved 3D consistent scenes at an infinite scale through unbounded layout extrapolation. Another set of works [3, 13, 41] focuses on indoor scene synthesis using costly 3D datasets [12] or CAD object retrieval [14, 15].
@@ 3.2 BEV-Point Initialization
In a local patch of a bird’s-eye-view (BEV) map comprising a semantic map $\mathbf{S}$ and a height field $\mathbf{H}$, a collection of BEV points within this patch can be produced by extruding the pixels in the semantic map $\mathbf{S}$ according to the corresponding values in the height field $\mathbf{H}$. We further introduce the binary density map $\mathbf{D}$ to adjust the sampling density for different semantic categories. This is driven by the observation that certain categories exhibit simpler textures (e.g., roads, water areas), allowing for reduced density to manage computational costs, while other categories (e.g., building façades) possess intricate textures, necessitating a greater number of points for representation. The coordinates of the generated BEV points, denoted as $\mathbf{C}_{F}\in\mathbb{R}^{N_{\rm pt}\times 3}$, can be generated as
@@ 3.3 BEV-Point Feature Generation
The style look-up table $\mathcal{T}$ queries the style code $\mathbf{z}_{T}^{i}\sim\mathcal{N}(0,1)$ for the instance $l$. Specifically,
@@ 4.3 Implementation Details
Hyperparameters. The grid size $g$ is set to $0.01$ and $N_{\rm PE}$ is set to $64$. The feature channels $C_{Z}$, $C_{F_{S}}$, $C_{F_{P}}$, and $C_{F_{T}}$ are set to $256$, $61$, $1280$, and $256$, respectively. The Gaussian attributes $\mathbf{A}$ consist of RGB values only (i.e., $C_{A}=3$). Initially, the loss weights $\lambda_{\rm L1}$, $\lambda_{\rm VGG}$, and $\lambda_{\rm GAN}$ are set to 10, 10, and 0.5, respectively. Afterward, $\lambda_{\rm L1}$ and $\lambda_{\rm VGG}$ are reduced by half every 20,000 epochs.
@@ 4.5 Ablation Studies
Table 2: Effectiveness of BEV-Point Representation. Note that “RayInt.” denotes “Ray Intersection”, respectively. The VRAM usage and inference time are measured on an NVIDIA Tesla A100.
Visibility Density #Points VRAM (G) Time (s)
31M Out of Memory
20M Out of Memory
Region 1.4M Out of Memory
Instance 680K 10.34 0.16
RayInt. 38.0K 1.41 0.10
RayInt. 31.7K 1.39 0.09
Table 3: Effectiveness of BEV-Point Decoder. Note that “Ser.” and “Pt.Trans.” denote “Point Serializer” and “Point Transformer”, respectively.
Ser. Pt.Trans. FID $\downarrow$ KID $\downarrow$ DE $\downarrow$ CE $\downarrow$
@@
Table 4: Comparison of different serialization methods. Note that “Eq. 13” denote the serialization used in Point Serializer.
@@
Effectiveness of BEV-Point Representation. The proposed BEV-Points, a compact scene representation, are essential for enabling 3D-GS to generate unbounded scenes by 1) considering point visibility during generation and 2) adjusting sampling density for different semantic categories using a binary density map. Table 2 compares different methods for considering point visibility, including no consideration, region-based, instance-based, and ray-intersection-based strategies for point selection. All methods keep the number of points constant: “region-based” selects points within the camera coverage, “Instance-based” selects visible instances, and “ray-intersection-based” identifies visible points through ray intersection [2]. The "ray-intersection-based" method is the most efficient compared to "region-based" and "instance-based" approaches, eliminating the most invisible points and reducing the total number of points by 97.7% and 96%, respectively. The density map is crucial in the BEV-Point representation because it reduces the number of points for semantic categories that have less texture and need fewer points. As shown in Table 2, removing the density map increases the point count by 20%-50%. As shown in Figure 1(b-c), the proposed BEV-Point representation maintains constant VRAM usage during training, unlike the original 3D-GS [23], which is impractical for unbounded scene generation due to excessive VRAM consumption. Additionally, it significantly reduces file storage as the number of points increases.
Effectiveness of BEV-Point Decoder. BEV-Point Decoder consists of two key components: Point Serializer and Point Transformer. As shown in Table 3 and Figure 6, removing both components leads to a significant degeneration due to the absence of spatial correlations. Introducing only the Point Transformer partially establishes spatial correlation, but it does not achieve the same effectiveness as using both components together.
Effectiveness of Serialization Methods. To investigate the impact of different serialization methods in point serializer, we compare the effects of using Equation 13, Hilbert curve [19] adopted in [8, 53], and both simultaneously on the generation results. As shown in Table 4, different serialization methods have minimal impact on the final generation results. We employ Equation 13 in Point Serializer because of its reduced computational complexity. Kindly refer to Figure 7 for more qualitative comparisons.
4.6 Limitations
While our method demonstrates promising results for unconditional 3D city generation, it still has several limitations. Firstly, the BEV-Point Initialization elevates points to a height based on the corresponding value in the height field, which assumes buildings adhere to the Manhattan assumption and cannot represent hollow structures. Secondly, the BEV-Point decoder does not fully exploit the expressive capacity of 3D-GS, as it only predicts RGB in the Gaussian attributes. However, predicting an excessive number of attributes simultaneously may destabilize the training process. Kindly refer to the discussion in Section A.2.
@@ 5 Conclusion
In this paper, we introduce GaussianCity for generating unbounded 3D cities. It is the first method to employ 3D-GS for unbounded scene generation. We introduce BEV-Point, a highly compact scene representation to mitigate the substantial increase in VRAM usage associated with representing large-scale scenes using 3D-GS. Additionally, we present BEV-Point Decoder, which leverages Point Serializer to capture the structural and contextual characteristics of unordered BEV points. Compared to previous methods, GaussianCity achieves better visual results and a 60-fold increase in runtime speed compared to CityDreamer. Extensive experimental results demonstrate the effectiveness of the proposed GaussianCity in generating both drone-view and street-view in city scenes.
@@ A.2 Discussion on Generated Gaussian Attributes
Due to the carefully designed BEV-Point initialization, all points are evenly distributed on the surface of objects. Therefore, Gaussian attributes other than RGB can be left unestimated and set to their default values. Table 5 and Figure 8 provide a quantitative and qualitative comparison of various attributes generated for 3D Gaussians, demonstrating that optimizing 3D Gaussian attributes beyond RGB not only complicates network convergence but also significantly impacts the quality of the generated results.
@@ B.1 Building Interpolation
As shown in Figure 9, GaussianCity showcases the capability to interpolate building styles controlled by the variable $z$. In the first row, local editing is applied by altering only the style code $z$ of the buildings within the bounding boxes, leaving the styles of the other buildings unchanged. In the second row, the style codes of all the buildings are interpolated from left to right, resulting in a gradual style transition across the entire row.
Figure 9: Linear interpolation along the building style. The style of each building gradually changes from left to right. In the first row, only the styles of the buildings enclosed by the bounding boxes are modified, while in the second row, the styles of all the buildings are changed.
@@ B.3 Additional Qualitative Comparison
In Figures 1112 and 13, we provide more visual comparisons with state-of-the-art methods on GoogleEarth and KITTI-360, respectively.